The time period of a gas bubble formed under water oscillating with a time period depending on static pressure P,density of water R and S total energy of explosion . Find relation between T,P

Grade:

2 Answers

Aman Bansal
11 years ago

Dear Pranamika,

Given T = k pa db Ec, where k is a proportionally constant. Substituting the dimensions of T, p, d and E, we have

      (T) = (ML-1T-2a (ML-3b (ML2-2c

Equating power of M, L, and T on both sides, we have

                         a + b + c = 0
                           –a –3b + 2c = 0
                          –2a – 2c = 1

Solving these equations, we get a = -5/6, b = 1/2 and c = 1/3

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. 

http://www.askiitians.com/packages/packages.aspx

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

Eeshnoor kaur
5 years ago
The time period of a gas bubble formed under water oscillating with a time period depending on static pressure P,density of water R and S total energy of explosion . Find relation between T,P